A random variable X has Poisson distribution with mean 2. Then P ( x > 1.5 ) equals____________.

Grade:11

2 Answers

Aman Bansal
11 years ago

Dear Shiva,

he formula for a poisson is P(X = x) = e^(-lambda) lambda ^(x) / x !

lambda = mean
also with a) do mean 3 <= x <= 5,

as this is not what you wrote, assuming that what I wrote is correct then just up the values of the formula from 3 to 5, for a

P(X = 3) + P(X = 4) + P(X = 5)

i.e. e^(-4) 4 ^(3) / 3 ! + e^(-4) 4 ^(4) / 4 ! + e^(-4) 4 ^(5) / 5 !

in b, you need to obtain 1 - P (x < 3), since P(x >= 3) would require you to add up every value from 3 to infinity, which would take forever, follow the example that I provided in a) to obtain P (x < 3), remember also what the minimum possible value of X is, (i.e. 0)

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

http://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

upender surepally
11 years ago

1-(3/e2)